# Dynamic Decoupling in Multidimensional Screening

### v1

@@ Abstract
I study multidimensional sequential screening. A monopolist contracts with an agent endowed with private information about the distribution of their eventual valuations of different goods; a contract is written and the agent reports their initial private information before drawing and reporting their valuations. In these settings, the monopolist frontloads surplus extraction: Any information rents given to the agent to elicit their post-contractual valuations can be extracted in expectation before valuations are drawn. This significantly simplifies the multidimensional screening problem. If the agent’s valuations satisfy invariant dependencies (valuations can be dependent across dimensions, but how valuations are coupled cannot vary in their initial private information), the optimal mechanism coincides with independently offering the optimal sequential screening mechanism for each good, regardless of the dependency structure.
Keywords: Dynamic Contracting, Dynamic Irrelevance, Multidimensional Screening.
@@ 1 Introduction
Multidimensional screening and dynamic contracting problems are generally difficult to solve: A multi-good monopolist’s revenue-maximizing mechanism often features nonlinear bundling and uncountably large menus. In this light, why does Amazon not bundle products? The answer is Amazon Prime: Extracting surplus intertemporally is more effective than at the moment of sale and simplifies the principal’s problem.
I study a model of dynamic multidimensional screening where a monopolist contracts with an agent who arrives with (one-dimensional) private information about their eventual distribution over (multidimensional) valuations; the model is a straightforward extension of CourtyLi2000 to multiple goods. Endowing the principal with the ability to contract intertemporally smooths out interdimensional incentive constraints, leading to separately selling different goods being optimal in a large variety of settings. I formalize a new force in dynamic adverse selection problems where the agent’s information arrives over time: The principal prefers to frontload surplus extraction, analogous to how transfers are backloaded in dynamic moral hazard problems. There are two forces, one governing feasibility and one governing optimality, for why surplus extraction is frontloaded. First is feasibility, coming from a pushforward force [^2] See, for example, the pullback lemma of luo2025payingpersuading.: Before the agent’s valuations are realized, the principal can extract the expected information rents the agent receives after their valuations are realized. Second is optimality. The principal needs to provide less information rents to the agent when the agent simply has less private information: Joint deviations have less bite as current period deviations can only be coordinated with future deviations in expectation. As such, the only information rents that need to actually be paid to the agent are those needed to elicit the agent’s pre-contractual information, in line with the dynamic irrelevance results of EsoSzentes2017. Post-contractual information rents can be provided “for free.”
Instead of providing the agent with their full static information rents, pulling surplus extraction forward distorts the agent’s information rents by (1) the inverse hazard ratio for their pre-contractual information and (2) the score of a likelihood ratio test for their valuations being drawn from the distribution corresponding to their report. This distortion not only works in favor of the principal, but additionally significantly simplifies the screening problem when the joint distribution of pre-contractual information and valuations satisfies invariant dependencies, generalizing conditional independence. Different dimensions of the agent’s valuations can be dependent, but the structure of these dependencies cannot vary in the agent’s pre-contractual private information. When this is the case, a likelihood ratio test is separable in each dimension and selling each good separately is optimal. The optimal contract has the same structure as CourtyLi2000: The agent is initially offered a menu of upfront payments and option payments for each good to be exercised when their valuation is drawn in the future. This mirrors real-world memberships at large retailers, such as Amazon Prime.
Finally, I compare my model and notion of dynamic irrelevance to other models of dynamic contracting. First, in an alternative setting where the agent once again starts with the same pre-contractual private information but valuations are revealed and goods are sold in sequence instead of simultaneously (as in segal2014), if valuations are independent then the monopolist’s optimal mechanism and profit are the same as the baseline case. Second, in the relaxed problem where an orthogonalized signal of the agent’s valuations were publicly observed (as in EsoSzentes2017) the monopolist’s optimal mechanism and profit are once again unchanged.
@@ 2 Model
The model is a straightforward extension of CourtyLi2000 to a multidimensional setting. A monopolist contracts with a single buyer. The monopolist sells $n$ goods. Before contracting, the buyer observes $\gamma\in[\underline{\gamma},\overline{\gamma}]=\Gamma$ drawn according to $G(\gamma)$. After contracting, the buyer draws $\theta\in\Theta=\prod_{j}[\underline{\theta}^{j},\overline{\theta}^{j}]\subseteq\mathbb{R}^{n}$ according to $F(\theta|\gamma)$. I assume $F,G$ have densities $f,g$ and $f$ is continuously differentiable with respect to $\gamma$ and $\theta$. Notationally, let superscripts denote marginals and subscripts denote partial derivatives, so $F^{j}_{\gamma}$ would denote the partial derivative of the marginal of $F$ in the $j$th dimension with respect to $\gamma$. For regularity, suppose the usual one-dimensional regularity condition holds marginal-by-marginal: for each marginal $j$, we have that $F^{j}_{\gamma}(\theta^{j}|\gamma)\leq 0$ and “virtual value”
is increasing in $\gamma$ and single-crossing in $\theta^{j}$.
@@
The constraints are as follows. First, $IC1$ pins down the buyer having no incentive to misreport $\gamma$ for any possible continuation misreport of $\theta$ after their new information arrives. However, their value of $\gamma$ still pins down the distribution of $\theta$ they face tomorrow. Next, $IC2$. After $\gamma$ is (truthfully) revealed, the buyer has no incentive to misreport $\theta$. Finally, $IR$ is an interim participation constraint: The buyer’s choice of whether or not to participate in the contract can depend on their observed value of $\gamma$.
3 Dynamic (Ir)relevance
We first present a rewriting of the problem that simplifies the monopolist’s problem and illustrates the role of frontloading surplus extraction.

### v2

@@ Abstract
I study multidimensional sequential screening. A monopolist contracts with a buyer who privately observes information about the distribution of their eventual valuations for multiple goods. After initial private information is reported and the contract is signed, the buyer learns and reports realized valuations. In these settings, the monopolist frontloads surplus extraction: Any information rents given to the buyer to elicit their true valuations can be extracted in expectation before those valuations are drawn, transforming the multidimensional screening problem by distorting buyer information rents compared to static screening. If the buyer’s distributions over valuations are commonly FOSD ordered and regular for each good; and satisfy invariant dependencies (valuations can be dependent across goods, but how valuations are coupled cannot vary), the optimal mechanism coincides with independently offering the optimal sequential screening mechanism for each good. This rationalizes membership payments followed by separate sales schemes seen across multiple industries.
Keywords: Dynamic Contracting, Multidimensional Screening.
@@ 1 Introduction
Firms often contract with consumers across multiple products and across multiple time periods. How do these distinct sources of multidimensionality—across space and time—interact and influence the design of optimal mechanisms? One strategy often used in practice is options-style pricing, where consumers are given the option to pay some upfront fee to receive lower prices later on. This scheme is seen in multiple industries. In retail, consider Costco’s “Gold Star” versus “Executive” memberships, Sam’s Club “Club” versus “Plus” memberships, or Amazon’s “standard” versus “Prime” accounts. In each case, the higher tier comes at a higher upfront payment but lower eventual costs in the form of cash-back on purchases, monthly credits, or free shipping. In air transportation (the classic application of sequential screening as in Courty and Li (2000)), many airlines offer the option to pay a higher upfront price to make flights refundable. In cloud computing, Google Cloud, for example, offers lower rates when committing to one or three years of usage compared to spot prices for compute.
Another feature of such pricing schemes is that the final prices offered to consumers are independent across goods. The price of eggs does not depend on milk purchases; the price of flights to Athens does not depend on planned trips to Barcelona; the price of M2 memory-optimized machine usage does not depend on C2 standard machine usage. This stands in sharp contrast to the multidimensional screening literature, which finds bundling (often in complicated and intractable ways) to be optimal. Why is it, then, that these products are not bundled?
To answer these questions, I study a model of dynamic multidimensional screening where a monopolist contracts with a buyer who arrives with private information about their distribution over eventual (multidimensional) valuations in period one when the contract is written. In period two, actual valuations are realized, goods are allocated, and transfers are made. The model is a straightforward extension of the canonical sequential screening setting introduced by Courty and Li (2000) to multiple goods. To build intuition behind separate sales, consider why bundling may occur in the multidimensional portion of this screening problem. In general, bundling goods helps the monopolist manage incentive constraints and reduce information rents given to the buyer, thereby capturing a larger share of total surplus generated from trade. However, endowing the monopolist with the ability to screen through time provides them with an alternative tool to reduce the effective information rents given to the buyer to elicit their eventual valuations: Any such information rents can be extracted in expectation before the buyer’s valuations are realized. As such, the only information rents that the buyer can actually capture are those for their initial private information (formalized in Lemma 1). This frontloading of surplus extraction is not only feasible per the previous discussion, but also optimal. The principal needs to provide less information rents to the buyer when the buyer simply has less private information: Joint deviations have less bite as current period deviations can only be coordinated with future deviations in expectation. Frontloading surplus extraction is similar to how transfers are often backloaded in dynamic moral hazard problems (see Ray (2002) or Luo (2025)).
Section 1.1 works through a two good example, with perfectly anti-correlated eventual values, and compares the optimal mechanism to the optimal pure bundling mechanism. Even in this case, where pure bundling completely negates any information rents the monopolist needs to pay to elicit the buyer’s final valuations, there are still large gains from using the optimal mechanism which features separate sales for either good. Bundling the two goods forces the buyer to potentially consume a good that they do not (really) want.
Several assumptions on the underlying environment are needed for a clean solution. First, the buyer’s initial private information can be ordered with, and thus parameterized by, a single dimension, with higher values corresponding to a higher (in the first-order stochastic dominance sense) distribution of eventual valuations for each good. While this does collapse the initial problem down to a one-dimensional screening problem, the space of outcomes is the set of all multidimensional screening mechanisms mapping realized valuations into allocations and transfers. As such, ex-ante without the decomposition developed in this paper, the problem remains multidimensional. Next, the joint distribution of initial information and eventual valuations must satisfy a sequential analogue of regularity for each good, commonly assumed in the literature. Finally, while final valuations may be coupled or correlated in any manner, this dependency structure cannot vary as the buyer’s initial private information varies. This new assumption, which I call invariant dependencies, arises only in the multidimensional case. When these conditions hold, the optimal mechanism takes a simple form. The optimal mechanism for each good on its own is to charge the buyer some upfront price to secure some eventual strike price, where buyers with higher initial private information being charged more upfront but less at the time of purchase. The optimal mechanism overall is to simply offer them the optimal individual-good mechanism separately: Charge the buyer the sum of each good’s upfront price to secure the same eventual strike prices for each good.
Absent invariant dependencies, (nonlinear) bundling may once again be part of the monopolist’s optimal mechanism. In this case, the monopolist can use heterogeneity in correlational structures as a tool to screen for valuations itself. Such screening is seen in practice as well: Consider vacations on cruise ships. Cruise lines offer different goods at different qualities such as better cabins, on-board internet, alcoholic beverages, specialty dining, and onshore excursions. When buying a low or medium-tier cabin, there are separate upgrade prices to purchase the aforementioned add-ons. Furthermore, these upgrade prices are independent of cabin quality; whether or not the cabin has a balcony is priced separately from whether or not it comes with internet. However, for top-tier cabins (the “family villa” on Norwegian Cruise Lines), each of the add-ons are bundled with the cabin purchase itself. This is rationalized by a situation where consumers with lower values of pre-contractual information have valuations for add-ons that are independent of one another (valuations for add-ons are essentially noise), but consumers with higher values of initial information have correlated valuations for add-ons (some premia for a vacation being “all-inclusive”).
The remainder of the paper proceeds as follows. Section 1.1 works through a simple example and compares pure bundling with the optimal mechanism. Section 2 discusses related literature while Section 3 sets up the model. Section 4 solves for the optimal mechanism and discusses how each assumption comes into play along. Section 5 compares the model in this paper to alternative models of multidimensional sequential screening to illustrate the strength of frontloading surplus extraction: The problem can be relaxed in several ways favoring the monopolist yet the firm attains as much revenue in the current model as any of the relaxations. Finally, Section 6 concludes.
1.1 Illustrative Example
Suppose a buyer is considering purchasing flights to Athens (Good $A$) or Barcelona (Good $B$). Their utility from $A$ is $v+\varepsilon$ and their utility from $B$ is $v-\varepsilon$ where $v$ is some initial private information (their demand for some sort of vacation) and $\varepsilon$ is some later shock (how alluring Athens is compared to Barcelona, or vice versa; the two may be substitutes if they only have a fixed number of vacation days at their non-tenure job). Initially, only $v$ is (privately) known to the buyer; $\varepsilon$ is (privately) revealed later on. Suppose $v$ is uniform over $[0,1/2]$ and $\varepsilon$ is uniform over [1/2,1/2][-1/2,1/2]. The seller writes a direct revelation mechanism contracting on both $v$ and $\varepsilon$.
Note that this specification stacks the deck in favor of bundling. Since the value of $A$ and $B$ are perfectly negatively correlated, the value of the bundle will always be $2v$ regardless of $\varepsilon$. By only selling the grand bundle, the seller does not need to give the buyer any information rents to elicit $\varepsilon$ and only needs to incentivize truthful elicitation of $v$. Under this regime, the problem turns into a one-dimensional screening problem selling to a buyer whose valuation of the bundle is uniform over $[0,1]$. As such, the optimal mechanism is to post a price of $1/2$ for the grand bundle. Producer surplus is $1/4$, consumer surplus is $1/8$, and total surplus is $3/8$. Bundling obviates the need to pay the buyer information rents to elicit $\varepsilon$ but comes at the cost of selling a good that the consumer does not want. Can the seller do better?
Instead of bundling, the seller can exploit the dynamic nature of the buyer’s private information to reduce buyer information rents. Suppose the seller no longer bundled the two goods so information rents for $\varepsilon$ still had to be paid. The seller knows the eventual information rents they need to give the buyer to elicit $\varepsilon$. As such, the seller can simply extract expected information rents before the buyer knows what $\varepsilon$ is. By doing this, the seller no longer needs to bundle; this avoids the issue of selling the buyer goods they do not want.
Concretely, consider the following mechanism (which we will later show to be optimal). The buyer reports $v$ and pays $2v^{2}$ upfront. Then, $\varepsilon$ is drawn and the seller posts a price of 1/2v1/2-v for each good independently (so the buyer can purchase neither good and pay zero, either one of the goods and pay 1/2v1/2-v, or both goods and pay 12v1-2v). Incentive compatibility for eliciting final valuations $v\pm\varepsilon$ comes for free since the second-stage mechanism is a posted price mechanism, which is incentive-compatible. Thus, it is sufficient to verify incentive-compatibility to truthfully report $v$. Suppose a buyer of type $v$ reports $\hat{v}$. They pay $2\hat{v}^{2}$ upfront and face a strike price of $1/2-\hat{v}$. The buyer thus purchases the first good if
$v+\varepsilon\geq\frac{1}{2}-\hat{v}\implies\varepsilon\geq\frac{1}{2}-\hat{v}-v$
so their expected surplus from the first good is
$\int_{\frac{1}{2}-\hat{v}-v}^{\frac{1}{2}}(v+\varepsilon-(1/2-\hat{v}))d\varepsilon=\left[(v+\hat{v}-1/2)\varepsilon+\frac{\varepsilon^{2}}{2}\right]_{\frac{1}{2}-\hat{v}-v}^{1/2}=\frac{1}{2}(v+\hat{v})^{2}.$
By symmetry, this is also their expected surplus from the second good; as such the interim utility of a buyer with type $v$ reporting $\hat{v}$ is
$\underbrace{\frac{1}{2}(v+\hat{v})^{2}}_{\text{Surplus from Good 1}}+\underbrace{\frac{1}{2}(v+\hat{v})^{2}}_{\text{Surplus from Good 2}}-\underbrace{2\hat{v}^{2}}_{\text{Upfront Payment}}=(v+\hat{v})^{2}-2\hat{v}^{2}.$
The first order condition with respect to $\hat{v}$ is
$2(v+\hat{v})-4\hat{v}=0\implies\hat{v}=v$
and the second order condition is satisfied as 2<0-2<0. Thus incentive compatibility is satisfied.
What are producer and consumer surplus from this mechanism? The seller gets $2v^{2}$ upfront and sells at each good at price of 1/2v1/2-v with probability $2v$ to a buyer of type $v$ and makes interim profits of
$2v^{2}+2(1/2-v)(2v)=2v-2v^{2}$
to a type $v$ buyer. Thus, integrating over $v$ gives expected producer surplus to be
$\int_{0}^{1/2}(2v-2v^{2})2dv=\left[2v^{2}-\frac{4v^{3}}{3}\right]_{0}^{1/2}=1/3$
which is higher than the surplus of $1/4$ from the pure bundling mechanism. Interim surplus of a type $v$ buyer is $4v^{2}-2v^{2}=2v^{2}$ (by plugging in $\hat{v}=v$ into the expression for the surplus a buyer of type $v$ gets when reporting $\hat{v}$). Thus, expected consumer surplus is
$\int_{0}^{1/2}(2v^{2})2dv=1/6$
which is higher than the surplus of $1/8$ from the pure bundling mechanism. As such, this alternative mechanism increases both producer and consumer surplus; total surplus increases from $1/4+1/8=3/8$ to $1/3+1/6=1/2$.
2 Related Literature
I relate to several strands of literature starting with the classic (separate) literatures on intertemporal screening and multidimensional screening. I also relate to recent work investigating the interaction between multidimensional screening and buyer information.
The sequential screening literature, started by Courty and Li (2000), considers a monopolist selling a single good to a buyer whose information arrives over time. Eső and Szentes (2017) generalizes the model to include buyer hidden actions and arbitrary time horizons and shows that the principle does not need to pay rents for any information that arrives after the buyer enters the contract; it is as if the principle observes the information for free as long as the desired mechanism is implementable. I apply this insight to the case of two periods but arbitrarily many goods in the second period. The multidimensional problem is more difficult than the multiple-period problem since the buyer in my problem can coordinate deviations across different goods whereas the buyer in Eső and Szentes (2007) cannot base current deviations on future information. See Eső and Szentes (2007), Pavan et al. (2014), Krähmer and Strausz (2017), Krasikov and Lamba (2022), and Li and Shi (2024) for additional work on sequential screening.
Next is the literature on multidimensional bundling, starting with Stigler (1963), Adams and Yellen (1976), McAfee et al. (1989), Armstrong (1996), Rochet and Choné (1998). Even when values are additive, the optimal mechanism features non-linear bundling and is generally intractable to find analytically (McAfee and McMillan (1988), Manelli and Vincent (2006), Pavlov (2011), Daskalakis et al. (2017a), Bergemann et al. (2022), Rochet and Stole (2003), Hart and Reny (2015), Hart and Reny (2019), Menicucci et al. (2015)). Another approach has been to start with a particular bundling strategy and then find conditions on the underlying problem for which that strategy is optimal (Haghpanah and Hartline (2021), Yang (2025)). This paper is particularly connected to Yang (2025), which also sets up the multidimensional screening problem with consumer heterogeneity ordered in a single dimension. After the monopolist collapses the multidimensional screening problem into a one-dimensional one via frontloading surplus extraction, revenues from selling incremental bundles given a fixed realization of post-contractual information are (1) single-crossing and (2) independent of other goods. As such, the optimal mechanism should exhibit nested bundling; in this paper, the system of upfront and seperate strike prices still induces nested bundling. Conditional on having the same realized final valuations, buyers that thought their distribution of valuations were higher pay more upfront and face lower eventual prices and thus choose to buy larger bundles at the prices they face.
Separate sales emerges in a multidimensional screening problem when robustness is a concern, developed in Carroll (2017) (and generalized in Che and Zhong (2025)). In Carroll (2017), the monopolist sells multiple goods and knows the buyer’s marginal distribution of values for each good but would like to maximize the worst-case guarantee against an adversarially chosen correlational structure. In this setting, selling each good independently is optimal as the monopolist cannot exploit interdependencies between goods due to the nature of the robustness exercise: Bundled sales perform worse at some potential joint distribution. In this paper, instead of being unable to exploit interdependencies, the monopolist does not need to exploit interdependencies, as frontloading surplus extraction is simply more effective.
Finally, there have been several recent papers combining learning (over time) and bundling. Lu and Zhao (2025) considers a monopolist selling two distinct goods, one in each of two distinct time periods. The buyer’s initial information influences both their valuation of the first good and the distribution of values of the second good. If higher first-good valuations lead to a first-order stochastic higher distribution of second-good valuations then the optimal mechanism consists of the optimal static screening mechanism for the first good and the optimal sequential screening mechanism for the second good. However, if higher first-good valuations lead to a first-order stochastic lower distribution of second-good valuations then the type for which individual rationality binds is endogenous and the optimal mechanism once again features bundling. Working with a single good in each period allows Lu and Zhao (2025) to abstract away from the cyclic monotonicity constraints that arise in multidimensional screening; this paper does not have any sales in the first period but an arbitrary number of goods in the second period. Individual rationality constraints being endogenous is why the distribution of valuations for each good must be ordered by first-order stochastic dominance in the same way in my model: Without this assumption, individual rationality constraints become endogenous and different goods may be used to provide participation incentives, thus negating dynamic decoupling and making the problem much more difficult.
Next, Frick et al. (2026) considers a multi-good monopolist who observes informative signals (that is, the seller perfectly knows the buyer’s valuations as the number of signals goes to infinity) about a buyer’s valuations. In this setting, pure bundling converges (as the number of signals goes to infinity) to full surplus extraction just as fast as the optimal mechanism whereas independent sales converges strictly slower. This paper, in contrast, endogenizes monopolist learning: Instead of exogenously observing informative signals, the buyer disclosing their initial information must be incentivized and is part of the contract.
Lastly, Pernoud and Yang (2025) considers a simultaneous-move game where buyers choose to learn a one-dimensional signal about a multi-dimensional product while a monopolist chooses a selling mechanism. In equilibrium, buyers choose vertical learning (a signal that is positively correlated with each of the product’s dimensions) while the seller’s mechanism features nested bundling. In their model, the buyer’s final purchasing decision is solely based on their signal while their true valuations are not part of the contracting problem. As such, frontloading surplus extraction is a key feature of my model not present in Frick et al. (2026) or Pernoud and Yang (2025).
@@ 3 Model
The model is a straightforward extension of Courty and Li (2000) to a multidimensional setting. A monopolist contracts with a single buyer. The monopolist sells $n$ goods. Before contracting, the buyer observes $\gamma\in[\underline{\gamma},\overline{\gamma}]=\Gamma$ drawn according to $G(\gamma)$. After contracting, the buyer draws $\theta\in\Theta=\prod_{j}[\underline{\theta}^{j},\overline{\theta}^{j}]\subseteq\mathbb{R}^{n}$ according to $F(\theta|\gamma)$. I assume $F,G$ have strictly positive densities $f,g$ and $f$ is continuously differentiable with respect to $\gamma$ and $\theta$. Notationally, let superscripts denote marginals and subscripts denote partial derivatives, so $F^{j}_{\gamma}$ would denote the partial derivative of the marginal of $F$ in the $j$th dimension with respect to $\gamma$.
@@
The constraints are as follows. First, $IC1$ pins down the buyer having no incentive to misreport $\gamma$ for any possible continuation misreport of $\theta$ after their new information arrives. However, their value of $\gamma$ still pins down the subsequent distribution of $\theta$. Next, $IC2$. After $\gamma$ is (truthfully) revealed, the buyer has no incentive to misreport $\theta$. Finally, $IR$ is an interim participation constraint: The buyer’s choice of whether or not to participate in the contract can depend on their observed value of $\gamma$.
3.1 Sufficient Assumptions
The following three conditions on the underlying environment are sufficient for multidimensional sequential screening problems to have a simple solution:
  1. A1.
    Commonly-Ordered: For all goods $j$ and $\gamma,\theta^{j}$,
    $F^{j}_{\gamma}(\theta^{j}|\gamma)\leq 0.$
    That is, higher values of $\gamma$ lead to higher distributions of $\theta^{j}$ in the first-order stochastic dominance order for all goods $j$.
  2. A2.
    Sequential Regularity: For each good $j$,
    is weakly increasing in $\gamma$ and $\theta^{j}$. [^2] Virtual values increasing in $\theta^{j}$ can be replaced with the weaker assumption of single-crossing instead. That is, the sequential analogue of a buyer’s “virtual value” satisfies regularity. This is standard in the sequential screening literature (for example, see Eső and Szentes (2007) or Lu and Zhao (2025)).
  3. A3.
    Invariant Dependencies: The distribution of percentiles
    $(F^{1}(\theta^{1}|\gamma),F^{2}(\theta^{2}|\gamma),...,F^{n}(\theta^{n}|\gamma))$
    is independent of $\gamma$. That is, valuations for each good can be coupled in arbitrary ways, but the coupling structure cannot vary as the buyer’s initial private information varies.
None of the three assumptions are required to show that the buyer only needs to be provided information rents to elicit their value of $\gamma$. Assumption A3 then establishes that the monopolist’s objective can be written to be additively separable in different goods. However, this does not immediately imply that separate sales is optimal as the different goods are still linked via the monopolist’s choice of how to optimally provide incentives for the buyer to truthfully report $\gamma$. If, additionally, assumptions A1 and A2 hold, then the solution to the relaxed problem is implementable so global incentive constraints have no bite, obviating the previous concerns.
What do these assumptions imply? Using the orthogonalization approach of Eső and Szentes (2007) or Eső and Szentes (2017) makes the above three assumptions more easily interpretable. By Sklar’s Theorem, there exists a unique distribution $C:[0,1]^{n}\times\Gamma\to[0,1]$ such that
and letting $c$ be the density of $C$,
Letting $z^{j}=F^{j}(\theta^{j}|\gamma)$ be the percentile of the buyer’s valuation of the $j$th good and $v^{j}$ denote the mapping from percentiles and initial private information to eventual valuations, we have that $\theta^{j}=(F^{j})^{-1}(z^{j}|\gamma)=v^{j}(z^{j},\gamma)$. In this setting, the three sufficient assumptions can be stated as follows:
  • A1’.
    Commonly-Ordered’: For each good $j$ and fixed $z^{j}\in[0,1]$,
    $v^{j}(z^{j},\gamma)$
    is increasing in $\gamma$. That is, for a fixed percentile, valuations are increasing in initial private information for all goods.
  • A2’.
    Sequential Regularity’: Virtual value for good $j$ being increasing in $\gamma$ is equivalent to
    $v^{j}_{\gamma,\gamma}/v^{j}_{\gamma}\leq v^{j}_{\gamma,z^{j}}/v^{j}_{z^{j}}.$
    That is, conditional on the same eventual valuation for good $j$, the buyer’s initial private information $\gamma$ has a smaller marginal impact on the final valuation as $\gamma$ grows.
    Virtual value for good $j$ being increasing in $z^{j}$ is equivalent to
    $v^{j}_{\gamma,z^{j}}\leq 0.$
    That is, buyers with higher initial private information $\gamma$ are (weakly) less impacted by the downstream shock $z^{j}$.
  • A3’.
    Invariant Dependencies’: The functions $C(z^{1},...,z^{n}|\gamma),c(z^{1},...,z^{n}|\gamma)$ can be written as $C(z^{1},...,z^{n}),c(z^{1},...,z^{n})$. That is, the cumulative and probability distribution functions of the percentiles do not depend on $\gamma$.
The following two examples illustrates what falls under the scope of the above three assumptions.
Example 1: Additive Independent Shock. Suppose
$\theta=u(\gamma)+\varepsilon$
where $\gamma$ follows any distribution with (weakly) decreasing inverse hazard rate, $u:\Gamma\to\mathbf{R}^{n}$ is deterministic and $\varepsilon$ is a random variable with support $\mathbf{R}^{n}$. As long as $u^{j}(\gamma)$ is (weakly) increasing and (weakly) concave in $\gamma$ for each $j$ and $\varepsilon$ is independent of $\gamma$, then assumptions A1-A3 are satisfied:
  • A1’.
    Since $\varepsilon$ is independent of $\gamma$, the same quantile $z^{j}$ corresponds to the same realization of $\varepsilon^{j}$ for all $\gamma.$ Then, $v^{j}(z^{j},\gamma)$ being increasing in $\gamma$ for fixed $z^{j}$ is equivalent to $u^{j}(\gamma)+\varepsilon^{j}$ being increasing in $\gamma$ for fixed $\varepsilon^{j}$.
  • A2.
    Sequential virtual value for good $j$ is
    $\theta^{j}-u_{\gamma}^{j}(\gamma)\frac{1-G(\gamma)}{g(\gamma)}$
    which is clearly increasing in $\theta^{j}$. If $u$ is (weakly) concave and (weakly) increasing and $G$ has (weakly) decreasing inverse hazard rate, then sequential virtual value is increasing in $\gamma$ as well.
  • A3’.
    Clearly satisfied.
Noise does not have to be perfectly additively separable: It could be that $\theta=u(\gamma)+\varepsilon(\gamma)$. If the magnitude of $\varepsilon$ does not vary too much as $\gamma$ varies, assumptions A1 and A2 will continue to hold. As long as the distribution of percentiles of $\varepsilon$ does not vary at all as $\gamma$ varies, assumption A3 will continue to hold.
This parametrization of the model also demonstrates how multidimensional sequential screening contrasts with the usual multidimensional screening problem. Consider two extreme cases of when information arrives. In the first case, suppose all of the buyer’s information arrives in $\gamma$ so $\varepsilon=0$ and the buyer’s valuation is just given by $u(\gamma)$. If this is the case, the only restriction is that $u(\gamma)$ is increasing in each component. This immediately collapses the problem into a one-dimensional screening problem; in particular, the marginal gains from adding any good to a given bundle is independent of the original bundle. As such, the marginal revenue intuition of Yang (2025) applies and separately posting prices for each good induces an allocation that features nested bundling, as buyers with higher values of $\gamma$ purchases larger (in the set order) bundles. In the second case, suppose all of the buyer’s information arrives in $\varepsilon$ so $u(\gamma)$ is constant in $\gamma$. In this case, the buyer has no (payoff-relevant) private information prior to contracting and the seller can achieve full surplus extraction by charging
$\mathbb{E}_{\varepsilon}\left[\sum_{j}\max\left\{u^{j}(\gamma)+\varepsilon^{j},0\right\}\right]$
upfront and posting a price for zero for each good.
Example 2: Noisy Vertical Learning. Suppose $\theta$ is normally distributed with mean $\mu$ and covariance matrix $\Sigma$. The buyer learns $\gamma=\alpha\cdot\theta+\varepsilon$ where $\varepsilon$ is normally distributed with mean 0 and variance $\sigma$. Suppose, similar to Pernoud and Yang (2025), [^3] Pernoud and Yang (2025) do not consider noise in the buyer’s signal but allow for elliptical distributions, which generalize normal distributions. Elliptical distributions would violate Invariant Dependencies as updating on $\gamma$ would then change the correlational structure. that learning is vertical; that is $\gamma$ is positively correlated with each dimension $\theta^{j}$. [^4] This implicitly puts assumptions on $\Sigma$; one sufficient condition used in Pernoud and Yang (2025) is to assume that all covariances are the same. Another possibility is to assume that all covariances are weakly positive; in general there cannot be too many covariances that are too negative. Then, the distribution of $\theta$ conditional on $\gamma$ is normal with mean
$\mu^{*}=\mu+\frac{\text{Cov}(\theta,\gamma)}{\text{Var}(\gamma)}(\gamma-E[\gamma])=\mu+\frac{\Sigma\alpha}{\alpha^{T}\Sigma\alpha+\sigma^{2}}(\gamma-\alpha^{T}\mu)$
and covariance
$\Sigma^{*}=\Sigma-\frac{\text{Cov}(\theta,\gamma)\text{Cov}(\theta,\gamma)^{T}}{\text{Var}(\gamma)}=\Sigma-(\Sigma\alpha)\left(\alpha^{T}\Sigma\alpha+\sigma^{2}\right)^{-1}(\alpha^{T}\Sigma).$
All three sufficiency assumptions are satisfied: A3 holds as $\Sigma^{*}$ is constant in $\gamma$. A1 holds since $\Sigma^{*}$ is constant in $\gamma$ and vertical learning implies that $\mu^{*}$ is increasing in $\gamma$. Finally, A2 holds since sequential virtual value for good $j$ is
$\theta^{j}-{\mu^{*}}_{\gamma}^{j}(\gamma)\frac{1-G(\gamma)}{g(\gamma)}$
and $\mu^{*}$ is linearly increasing in $\gamma$ while the distribution $G$ has decreasing inverse hazard rate as $\gamma$ is normally distributed.
4 Dynamic Decoupling
The main analysis proceeds in two steps: (1) formalizing the intuition that the monopolist can frontload surplus extraction to re-write the monopolist’s problem and (2) showing that the resulting distortion to buyer information rents induces independent sales.
The first step: Frontloading surplus extraction. Let $u(\gamma,\theta)=\theta\cdot q(\gamma,\theta)-t_{2}(\gamma,\theta)$ be the buyer’s utility net of $t_{1}$. For any $u(\gamma,\theta)$, applying the envelope theorem to period two gives that the allocation rule must be $q(\gamma,\theta)=\nabla_{\theta}u(\gamma,\theta)$. As such, total surplus ($\theta\cdot\nabla_{\theta}u(\gamma,\theta))$ being equal to consumer surplus ($u(\gamma,\theta)$) plus producer surplus ($t_{2}(\gamma,\theta$) implies that period two transfers are the same as in standard multidimensional screening settings:
Next, let
$U(\gamma^{\prime}|\gamma)=\int_{\Theta}u(\gamma^{\prime},\theta)dF(\theta|\gamma)-t_{1}(\gamma^{\prime});U(\gamma)=U(\gamma|\gamma)=\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)-t_{1}(\gamma).$
be the interim utility a consumer with type $\gamma$ receives when reporting $\gamma^{\prime}$ and the interim utility of a type $\gamma$ consumer when reporting truthfully. Since $\gamma$ is not directly payoff-relevant and only affects payoffs by influencing the distribution of $\theta$, incentive compatibility of truthfully revealing $\theta$ holding for all $\gamma$ implies that even after a deviation to $\gamma^{\prime}$, it is still a best response to truthfully report $\theta$. Then, by incentive compatibility for revealing $\gamma$, we have $\gamma\in\operatorname*{\arg\!\max}_{\gamma^{\prime}}U(\gamma^{\prime}|\gamma)$ so the Envelope Theorem applies and local incentive constraints implies
By the Fundamental Theorem of Calculus, we have
As usual, we can set $U(\underline{\gamma},\underline{\gamma})=0$. Equating this with the direct computation of
$U(\gamma)=\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)-t_{1}(\gamma)$
as expected downstream utility minus transfers and solving for transfers gives
Thus, $t_{1},t_{2},q$ are all pinned down by $u$ so the monopolist only needs to maximize over $u$; the monopolist’s objective function can be written as
Observe that the first term of $\mathbb{E}_{\gamma}\left[t_{1}(\gamma)\right]$ and the second term of $\mathbb{E}_{\gamma,\theta}\left[t_{2}(\gamma,\theta)\right]$ cancel out: Any information rents given to the consumer for their post-contractual information can be extracted, in expectation, before that information arrives. The remainder of the computation is standard, interchanging the two integrals with respect to $\gamma$. Expected transfers are
$=$ $\mathbb{E}_{\gamma,\theta}\left[\theta\cdot\nabla_{\theta}u(\gamma,\theta)-u(\gamma,\theta)\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}\right].$
Putting the above computation together, monopolist’s problem is as follows: